OpenStudy (anonymous):

@satellite73, how do you reckon I should state the zeros....

OpenStudy (anonymous):

like this

OpenStudy (anonymous):

I dont understand the format it should be in..

OpenStudy (anonymous):

ok we factored it right?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

(3x-1)(3x+1)(3x-i)(3x+i)

OpenStudy (anonymous):

then we set each factor equal to zero and solve for x ( in our heads)

OpenStudy (anonymous):

which is positive and negative 1/3 but what about the ones in paranthesis with i?

OpenStudy (anonymous):

\[3x-1=0\implies x=\frac{1}{3}\] etc

OpenStudy (anonymous):

same with them \[3x-i=0\] \[x=\frac{i}{3}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

let me see if it accepts it

OpenStudy (anonymous):

you will have 4 zeros, each with multiplicity 1

OpenStudy (anonymous):

okay thank u!

OpenStudy (anonymous):

yw